OpenStudy (cresciez):

Trigonometry help? :) If the cos 30° = square root 3 over 2 , then the sin 60° = _____.

OpenStudy (cresciez):

@kayders1997 ? ^-^

OpenStudy (candy13106):

square root 3 over 2, because the angles are complementary sin(70)= cos(20) tan x = sin x / cos x z/10 = sin x / (10/y) sin x = (z/10)(10/y) = z over y

OpenStudy (cresciez):

I'm not sure I understand... I saw this on yahoo btw lol How do you get z and y?

OpenStudy (cresciez):

I know the answer but I need to be walked through it

OpenStudy (cresciez):

Thank you for trying though! :)

OpenStudy (cresciez):

@Michele_Laino maybe? c:

OpenStudy (cresciez):

@Nnesha is it possible you could help? ^^

Nnesha (nnesha):

sure. do you know the relationship of cos sin and theta ??

Nnesha (nnesha):

 \[\rm sin \rm \theta = \frac{ opposite }{ hypotenuse }~~~~ \cos \theta = \frac{ adjacent }{ hypotenuse } ~~\tan \theta = \frac{ opposite }{ adjacent }\]

Nnesha (nnesha):

use the given informaton cos 30 degree = sqrt{3}/2 and cos = adjacent over hypotenuse write these values on the right triangle  it's a right triangle so if one angle is 30 what would be the 3rd angle equal to ?? and how would you find 3rd side of right triangle ?any idea :=))